OpenStudy (anonymous):

Solve the following equations for x. Identify any extraneous solutions. Show all work. -4√x+9=20

OpenStudy (anonymous):

It isn't clear if the 9 is also under the radical..

OpenStudy (anonymous):

I disagree.

OpenStudy (anonymous):

\[-4\sqrt{x+9} = 20\]\[\implies \sqrt{x+9} = -5\] And the square root cannot be negative, so there are no solutions. And if the problem was: \[-4\sqrt{x} + 9 = 20\]\[\implies -4\sqrt{x} = 11\]\[\implies \sqrt{x} =-\frac{11}{4}\] Which also has no solutions.

OpenStudy (anonymous):

zero displacement

OpenStudy (anonymous):

x=16 is a valid solution. \[(-4)(-5)=20\]Or,\[-4\sqrt{25}=-4\sqrt{16+9}\]

OpenStudy (anonymous):

So lets see: \[-4\sqrt{x+9} = 20\]\[x = 16 \implies -4\sqrt{16 + 9} = 20\]\[\implies \sqrt{25} = -5\]\[\implies 5 = -5\] Yeah, nope.

OpenStudy (anonymous):

Even your own version isn't right because you are still equating -5 with sqrt(25). They aren't equal.

OpenStudy (anonymous):

\[\sqrt{25}=\pm5\] as in:\[\sqrt{(-5)^2}=\sqrt{5^2}=\sqrt{25}\]

OpenStudy (anonymous):

There is so much wrong with that...

OpenStudy (anonymous):

sqrt25 does indeed equal -5

OpenStudy (anonymous):

\[\sqrt{25} = 5\] Always.

OpenStudy (anonymous):

I will grant you that \[\sqrt{25} = \sqrt{(-5)^2}\]But that's because \[ \sqrt{(-5)^2} = 5\]

OpenStudy (anonymous):

You cannot take the square root of a number and get a negative.

OpenStudy (anonymous):

Yes, you can.

OpenStudy (anonymous):

Show me one number 'q' such that sqrt(q) < 0

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Square_root

OpenStudy (anonymous):

I know the definition of the square root. Did you come up with a q?

OpenStudy (anonymous):

Every positive real number has TWO square root; one positive and one negative. So yeah..q is true for all positive numbers on the real number line. Pick one.

OpenStudy (anonymous):

From the wikipedia article you quoted..: Every non-negative real number x has a unique non-negative square root, called the principal square root, denoted by a radical sign as \( \sqrt{x}\). So yes, there are postive and negative square roots, the \(\sqrt{x}\) is defined to be the positive one.

OpenStudy (anonymous):

So you can certainly say that the equation: \[x^2 = 3\]Has solutions \(x = \pm \sqrt{3}\) But there are no solutions for which \(\sqrt{x} = q\) such that \(q < 0\)

OpenStudy (anonymous):

Sorry had to clarify that the last equation wasn't related to the first two.

OpenStudy (anonymous):

The square root function is the inverse function of the exponential function. Starting with the exponential function, we have: \[f(x)=x^2\]This function has domain and range:\[x \in \mathbb{R} ; y \ge0\]Now we find its inverse in the usual way:\[x=y^2\]and,\[x=\sqrt{y} \]and,\[y=\sqrt{x}\]The domain and range of an inverse function are swapped with the original function. i.e., the domain the square root function is the range of the exponential function:\[x \ge0\]No biggie, we still agree on this. The range of the square root function is the domain of the exponential function:\[y \in \mathbb{R}\]This obviously includes negative solutions. The domain and range are swapped for inverse functions because an inverse functions "undo" one another.

OpenStudy (anonymous):

Good debate. I doubt my last one convinced you though. Off to bed..to be continued...

OpenStudy (anonymous):

Only one to one functions have inverse functions.

OpenStudy (anonymous):

This is why \(\sqrt{x} = q\) is defined to be the positive number q such that \(q^2 = x\).

OpenStudy (anonymous):

Otherwise one input (25 for example) would have many ouputs (5 and -5). Functions don't behave this way. For one input there is one and only one output. So the function \(\sqrt{x}\) is defined to yield the positive value only. If you want a negative, you put a negative out in front of it (\(-\sqrt{x}\)) Or if you want both you can have both. \(\pm\sqrt{x}\)

OpenStudy (anonymous):

It looks like I'm wrong on this one. http://www.wolframalpha.com/input/?i=-4 √%28x%2B9%29%3D20+